# EN-TensorCore: Advancing TensorCores Performance through Encoder-Based Methodology

### v1

@@ Abstract.
Tensor computations, with matrix multiplication being the primary operation, serve as the fundamental basis for data analysis, physics, machine learning, and deep learning. As the scale and complexity of data continue to grow rapidly, the demand for tensor computations has also increased significantly. To meet this demand, several research institutions have started developing dedicated hardware for tensor computations. To further improve the computational performance of tensor process units, we have reexamined the issue of computation reuse that was previously overlooked in existing architectures. As a result, we propose a novel EN-TensorCore architecture that can significantly reduce chip area and power consumption. Furthermore, our method is compatible with existing tensor processing architectures. We evaluated our method on prevalent microarchitectures, the results demonstrate an average improvement in area efficiency of 8.7%, 12.2%, and 11.0% for tensor computing units at computational scales of 256 GOPS, 1 TOPS, and 4 TOPS, respectively. Similarly, there were energy efficiency enhancements of 13.0%, 17.5%, and 15.5%.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06submissionid: 874
@@ 1. Introduction
The surge in data and intricacy has significantly boosted the need for tensor-based computations, as AI has evolved from a marginal area research topic to a core driver of modern technology over the last decade. AI applications, including search engines, generative AI based on Large Language Models (LLMs), and tools for self-driving cars, have become widespread in everyday life. Many of these AI applications rely on large-scale tensor computations, leading to specialized hardware development for these calculations by researchers. Include NVIDIA’s incorporation of Tensor Cores(nvidia, ) in their Volta, Turing, Ampere, and the latest Hopper architectures; Google’s Tensor Processing Unit (TPU)(googletpu, ); Graphcore’s MK series Intelligence Processing Unit (IPU)(ipu, ); SambaNova’s SN(sambanova, ) series deep learning processors; HUAWEI’s Ascend(ascend, ); and Cambricon’s Machine Learning Unit (MLU)(mlu, ). For mobile devices, the SnapDragon 8gen3(snapdragon, ), MediaTek Dimensity 9300(mediatek, ), and Apple A17(apple, ) have all integrated neural processing units to enhance the energy efficiency of generative AI in mobile SoCs.
@@ 3.1. EN-TensorCore Architecture
When we extend this behavior to arrays of multipliers, since matrix multiplication or convolution has phenomenon of same multiplicand by different multiplier in the spatial dimension or time dimension, there is a repeated encoding behavior of the multiplicand $A$ inside the multiplier of PEs (Fig. 3(a)). From the perspective of TCUs, what is needed is the encoded multiplicand $A$, not $A$ itself. When applied to the existing various TCUs hardware architectures, whether it’s based on data broadcasting like 2D Matrix or 1D/2D Array, or data flow-based like Systolic Array or 3D Cube, enhancing performance only requires the following simple steps: First, remove the encoder logic from all multipliers within the tensor cores (Fig. 3(b)), retaining only the partial product compressor and the full adder; second, add a single encoder to each column of the multiplicand pathway outside the array (Fig. 3(c)), allowing the encoded multiplicand $A$ to flow or broadcast within the array. In terms of area, this can reduce the size of an individual PE, and when applied to TCUs with large-scale multiplier arrays, it can make the array layout more efficient and compact, which is beneficial for reducing the latency of multiplication operations. In terms of power consumption, it not only saves the power consumption caused by repetitive encoding logic, but also, due to the reduced array area, makes the data transmission pathways between adjacent PEs shorter, which further reduces the power consumption associated with data flow between PEs.
@@ 3.2. Challenge in Modified Booth Encoding
However, applying MBE to the EN-TensorCore architecture does not achieve the expected effect. This is because every 2 bits of the multiplicand need to be encoded into 3 bits to serve as control lines NEG, SE, CE. Thus, for the multiplication of two $n$-bit numbers, the multiplicand needs to be encoded into $[n/2]*3$ bits. Externalizing the encoder would actually cause the width of the interconnects between multiplicands in the TCUs, which is undesirable. The increase in interconnect width significantly affects the chip’s area and power. To address this, we redesigned an encoder that encodes an $n$-bit multiplicand into $n+1$ bits to reduce the width of the encoded numbers, which will be described in the subsection 3.3.
3.3. Modified Encoding in EN-TensorCore
@@ 3.3.1. Construction of Encoding Polynomials
Based on the foregoing analysis, we have demonstrated that for $n=2$ and $n=m$, through the recursive expressions Eq. 16 and Eq. 17, a set of solutions $\{Cin_{N+1},w_{N},w_{N-1},...,w_{0}\}$ can be identified, enabling the polynomial $Q_{N}$ to be expressed as the polynomial $Q^{\prime}_{N}$. According to the principle of mathematical induction, for all $N\geq 2$, we can employ a recursive method to represent the polynomial $Q_{N}$ as the polynomial $Q^{\prime}_{N}$. Hence, based on Eq. 7, 8, 16 and 17, we can encode an $n$-bit unsigned number into $n/2$ two-bit coefficients and one one-bit coefficient.
For example, 78(int8) is encoded as $Encode(78)=\{0,1,1,-1,2\}$, therefore the multiplication result of $B$ and 78 is $B4^{3}+B4^{2}-B4^{1}+2B$. The subsequent calculations are completely identical to those of MBE, that is, $B$ with different bit weights is added together to obtain the final multiplication result. When $A$ is negative, it is only necessary to perform a transformation on $B$ based on the sign bit of $A$ to implement signed multiplication (at this point, the hardware will choose $-B$ as the $B$ in the above formula to participate in the actual calculation).
In terms of the number of encoders, for an $n$-bit signed number using MBE, $n/2$ encoders are required; we need $(n/2-1)$ encoders (as the lowest 2 bits do not need encoding). For the encoded bit width, MBE requires $[n/2]*3$ bits, whereas our method requires $n+1$ bits. In the experimental section, we will first perform an evaluation of the area, delay, and other performance metrics of encoders for different bit widths, and test the performance of the optimized encoders within the EN-TensorCore.
@@
Figure 5. Experiments of Area and Power of EN-TensorCore
4. Experiment
4.1. Implement Environment
We implement our design in RTL and then synthesize, place and route it with Synopsys Design Compiler toolchain using the SMIC 40nm NLL-HS-RVT technology. We evaluate the performance and energy costs using PrimeTime PX based on VCD waveform files obtained from simulation with typical corner process.
4.2. Performance of the Encoder
In Table. 1, we ran performance tests on encoders for both 2-bit and multi-bit cases. In the comparison of resource consumption for 2-bit encoders, our method requires one less AND gate but one additional XNOR gate compared to MBE. This is because the XOR logic is used to generate the sum for both 2 bits in Eq. 17. Although our individual encoder is at a disadvantage in terms of area, in the case of multi-bit encoding tests, our encoding representation requires one less encoder due to the fact that the lowest 2 bits do not need encoding. Therefore, our method only exhibits advantages in terms of area and power consumption when the encoding bit width is less than 14 bits. In terms of latency, due to the parallel computation nature of MBE encoding, the extended bit width has almost no impact on its latency. However, our encoding is based on carry-chain encoding, causing the latency to gradually increase with the width of the multiplicand. This is a drawback of our method, but it becomes an insignificant factor in the EN-TensorCore architecture, as the encoders will be placed outside the array and enter the array through registers. At this point, the true critical path lies within the adders and accumulators inside the PE. Lastly, in terms of the comparison of encoding bit width, our encoding representation is not sensitive to the data bit width, whereas MBE’s data bit width increases by a factor of 1.5. From the perspective of a single multiplier, this may not be an important factor. However, in the EN-TensorCore architecture, it becomes an obstacle that limits the performance improvement of the array. We also conducted performance tests on multipliers, using Synopsys DesignWare standard process library multiplier (DW IP) and Modified Booth Multiplier as benchmarks. In terms of INT8 performance, the area and power consumption are comparable to those of DW IP, with a slightly higher delay of 0.12ns. However, this is not the primary factor. We designed it specifically for the EN-TensorCore architecture. In the experiments after removing the encoder logic (RME_Ours), there are significant improvements in area, delay, and power consumption, making it a promising solution for achieving significant performance improvements in large-scale TCUs.
Figure 6. Up Ratio of Area Efficiency and Power Efficiency
4.3. Performance of EN-TensorCore
We did performance tests on EN-TensorCore within the scope of mainstream AI accelerators, covering various tensor computation microarchitectures: 2D Matrix (Fig. 2(a)), 1D/2D Array (Fig. 2(b)), two types of Systolic Array (WS and OS) (Fig. 2(c)(d)), and 3D Cube (Fig. 2(e)). The array sizes will be tested at $16^{2}$, $32^{2}$, $64^{2}$ (Cube: $4^{3}$, $8^{3}$, $16^{3}$) to evaluate the scalability, energy efficiency, and area efficiency of the EN-TensorCore architecture (Fig. 3(c)). The benchmark object for testing is the PE composed of Synopsys DesignWare standard process library multiplier. The EN-TensorCore architecture utilizes two encoders (MBE and Our Encoder) with register outputs, and all test on 500MHZ. The accuracy of multipliers in the PE is INT8 (Due to the prevalent use of INT8 in existing AI accelerators and edge devices, where the mantissa part of floating-point numbers often involves low-bit-width unsigned multiplication (BF16(UINT7), TF32(UINT10)), we adopt INT8 as the multiplication bit width for testing TCUs.), and the accumulator width is $16+\log_{2}S$. $S$ represents the array size.
[…]
@@ 4.3. Performance of EN-TensorCore
In the tests of energy efficiency and area efficiency (Fig. 6(a)(b)), we mainly compare the scalability of individual TCUs with different computational scales under EN-TensorCore. Due to the square relationship between the number of removed encoders and the array size, increasing the computational scale to a certain extent will result in higher performance. As shown in Fig. 6(a)(b), when the computational scale of our encoder-based EN-TensorCore expands from 256GOPS to 1TOPS and 4TOPS, the average area efficiency improves from 8.7% to 12.2% and 11.0%, and the energy efficiency improves from 13.0% to 17.5% and 15.5%, respectively. Among them, the 1D/2D Array achieves a 20.2% increase in area efficiency and a 20.5% increase in energy efficiency compared to the baseline at 1TOPS. This is due to the specific characteristics of the multiplier-adder architecture itself (with no PEs, multipliers and multiplicands are not pipelined to the adder tree). In this case, the performance improvement of the EN-TensorCore architecture is the highest.

### v2

# EN-T: Optimizing Tensor Computing Engines Performance via Encoder-Based Methodology

@@ Abstract.
Tensor computations, with matrix multiplication being the primary operation, serve as the fundamental basis for data analysis, physics, machine learning, and deep learning. As the scale and complexity of data continue to grow rapidly, the demand for tensor computations has also increased significantly. To meet this demand, several research institutions have started developing dedicated hardware for tensor computations. To further improve the computational performance of tensor process units, we have reexamined the issue of computation reuse that was previously overlooked in existing architectures. As a result, we propose a novel EN-T architecture that can reduce chip area and power consumption. Furthermore, our method is compatible with existing tensor processing units. We evaluated our method on prevalent microarchitectures, the results demonstrate an average improvement in area efficiency of 8.7%, 12.2%, and 11.0% for tensor computing units at computational scales of 256 GOPS, 1 TOPS, and 4 TOPS, respectively. Similarly, there were energy efficiency enhancements of 13.0%, 17.5%, and 15.5%.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
The surge in data and intricacy has significantly boosted the need for tensor-based computations, as AI has evolved from a marginal area research topic to a core driver of modern technology over the last decade.
@@
AI applications, including search engines, generative AI based on Large Language Models (LLMs), and tools for self-driving cars, have become widespread in everyday life. Many of these AI applications rely on large-scale tensor computations, leading to specialized hardware development for these calculations by researchers. Include NVIDIA’s incorporation of Tensor Cores(nvidia, ) in their Volta, Turing, Ampere, and the latest Hopper architectures; Google’s Tensor Processing Unit (TPU)(googletpu, ); Graphcore’s MK series Intelligence Processing Unit (IPU)(ipu, ); SambaNova’s SN(sambanova, ) series deep learning processors; HUAWEI’s Ascend(ascend, ); and Cambricon’s Machine Learning Unit (MLU)(mlu, ). For mobile devices, the SnapDragon 8gen3(snapdragon, ), MediaTek Dimensity 9300(mediatek, ), and Apple A17(apple, ) have all integrated neural processing units to enhance the energy efficiency of generative AI in mobile SoCs.
@@ 3.1. EN-T Architecture
When we extend this behavior to arrays of multipliers, since matrix multiplication or convolution has phenomenon of same multiplicand by different multiplier in the spatial dimension or time dimension, there is a repeated encoding behavior of the multiplicand $A$ inside the multiplier of PEs (Fig.3(a)). From the perspective of TCUs, what is needed is the encoded multiplicand $A$, not $A$ itself. When applied to the existing various TCUs hardware architectures, whether it’s based on data broadcasting like 2D Matrix or 1D/2D Array, or data flow-based like Systolic Array or 3D Cube, enhancing performance only requires the following simple steps: First, remove the encoder logic from all
Figure 4. Modified Booth Multiplier
multipliers within the tensor cores (Fig.3(b)), retaining only the partial product compressor and the full adder; second, add a single encoder to each column of the multiplicand pathway outside the array (Fig.3(c)), allowing the encoded multiplicand $A$ to flow or broadcast within the array. In terms of area, this can reduce the size of an individual PE, and when applied to TCUs with large-scale multiplier arrays, it can make the array layout more efficient and compact, which is beneficial for reducing the latency of multiplication operations. In terms of power consumption, it not only saves the power consumption caused by repetitive encoding logic, but also, due to the reduced array area, makes the data transmission pathways between adjacent PEs shorter, which further reduces the power consumption associated with data flow between PEs.
@@ 3.2. Challenge in Modified Booth Encoding
However, applying MBE to the EN-T architecture does not achieve the expected effect. This is because every 2 bits of the multiplicand need to be encoded into 3 bits to serve as control lines NEG, SE, CE. Thus, for the multiplication of two $n$-bit numbers, the multiplicand needs to be encoded into $[n/2]*3$ bits. Externalizing the encoder would actually cause the width of the interconnects between multiplicands in the TCUs, which is undesirable. The increase in interconnect width significantly affects the chip’s area and power. To address this, we redesigned an encoder that encodes an $n$-bit multiplicand into $n+1$ bits to reduce the width of the encoded numbers, which will be described in the subsection 3.3.
3.3. Modified Encoding in EN-T architecture
@@ 3.3.1. Construction of Encoding Polynomials
Figure 6. Experiments of Area and Power of EN-T architecture
Based on the foregoing analysis, we have demonstrated that for $n=2$ and $n=m$, through the recursive expressions Eq. 16 and Eq. 17, a set of solutions $\{Cin_{N+1},w_{N},w_{N-1},...,w_{0}\}$ can be identified, enabling the polynomial $Q_{N}$ to be expressed as the polynomial $Q^{\prime}_{N}$. According to the principle of mathematical induction, for all $N\geq 2$, we can employ a recursive method to represent the polynomial $Q_{N}$ as the polynomial $Q^{\prime}_{N}$. Hence, based on Eq. 7, 8, 16 and 17, we can encode an $n$-bit unsigned number into $n/2$ two-bit coefficients and one one-bit coefficient. Modified encoding logic as show in Fig.5.
For example, 78(int8) is encoded as $Encode(78)=\{0,1,1,-1,2\}$, where the first number is the sign, and the next four 2-bit numbers are the encoding numbers, therefore the multiplication result of $B$ and 78 is $B4^{3}+B4^{2}-B4^{1}+2B$. The subsequent calculations are completely identical to those of MBE, that is, $B$ with different bit weights is added together to obtain the final multiplication result. When $A$ is negative, it is only necessary to perform a transformation on $B$ based on the sign bit of $A$ to implement signed multiplication (at this point, the hardware will choose $-B$ as the $B$ in the above formula to participate in the actual calculation).
In terms of the number of encoders, for an $n$-bit signed number using MBE, $n/2$ encoders are required; we need $(n/2-1)$ encoders (as the lowest 2 bits do not need encoding). For the encoded bit width, MBE requires $[n/2]*3$ bits, whereas our method requires $n+1$ bits. In the experimental section, we will first perform an evaluation of the area, delay, and other performance metrics of encoders for different bit widths, and test the performance of the optimized encoders within the EN-T architecture.
4. Experiment
4.1. Implement Environment
We implement our design in RTL and then synthesize, place and route it with Synopsys Design Compiler toolchain using the SMIC 40nm NLL-HS-RVT technology and using ARM Memory Compiler to generate on-chip SRAM. We evaluate the performance and energy costs using PrimeTime PX based on VCD waveform files obtained from simulation with typical corner process.
@@ 4.1. Implement Environment
Table 1. Area/$\mu m^{2}$, Delay/$ns$ and Power/$\mu W$
Figure 7. Up Ratio of Area Efficiency and Power Efficiency
4.2. Performance of the Encoder
In Table. 1, we ran performance tests on encoders for both 2-bit and multi-bit cases. In the comparison of resource consumption for 2-bit encoders, our method requires one less AND gate but one additional XNOR gate compared to MBE. This is because the XOR logic is used to generate the sum for both 2 bits in Eq. 17. Although our individual encoder is at a disadvantage in terms of area, in the case of multi-bit encoding tests, our encoding representation requires one less encoder due to the fact that the lowest 2 bits do not need encoding. Therefore, our method only exhibits advantages in terms of area and power consumption when the encoding bit width is less than 14 bits. In terms of latency, due to the parallel computation nature of MBE encoding, the extended bit width has almost no impact on its latency. However, our encoding is based on carry-chain encoding, causing the latency to gradually increase with the width of the multiplicand. This is a drawback of our method, but it becomes an insignificant factor in the EN-T architecture, as the encoders will be placed outside the array and enter the array through registers. At this point, the true critical path lies within the adders and accumulators inside the PE. Lastly, in terms of the comparison of encoding bit width, our encoding representation is not sensitive to the data bit width, whereas MBE’s data bit width increases by a factor of 1.5. From the perspective of a single multiplier, this may not be an important factor. However, in the EN-T architecture, it becomes an obstacle that limits the performance improvement of the array. We also conducted performance tests on multipliers, using Synopsys DesignWare standard process library multiplier (DW IP) and Modified Booth Multiplier as benchmarks. In terms of INT8 performance, the area and power consumption are comparable to those of DW IP, with a slightly higher delay of 0.12ns. However, this is not the primary factor. We designed it specifically for the EN-T architecture. In the experiments after removing the encoder logic (RME_Ours), there are significant improvements in area, delay, and power consumption, making it a promising solution for achieving significant performance improvements in large-scale TCUs.
Global Buffer
Size($KB$) Area($\mu m^{2}$) Read Power($W$) Write Power($W$)
256 614400 0.0205 0.04515
Activation and Weight Buffer
Size($KB$) Area($\mu m^{2}$) Read Power($W$) Write Power($W$)
64 153600 0.0146 0.0322
SIMD Vector Processing Engine
ALU Precision Area($\mu m^{2}$) Power($W$)
32 TF32 126481 0.0951
Controller and Img2col
Number Area($\mu m^{2}$) Power($W$)
2 83679 0.0632
Encoder
Number Area($\mu m^{2}$) Power($W$)
32 1895.36 0.00089
Table 2. On-chip Parameters of the SoC Benchmark
Figure 8. Benchmark SoC
4.3. Performance of EN-T architecture
We did performance tests on EN-T architecture within the scope of mainstream AI accelerators, covering various tensor computation microarchitectures: 2D Matrix (Fig.2(a)), 1D/2D Array (Fig.2(b)), two types of Systolic Array (WS and OS) (Fig.2(c)(d)), and 3D Cube (Fig.2(e)). The array sizes will be tested at $16^{2}$, $32^{2}$, $64^{2}$ (Cube: $4^{3}$, $8^{3}$, $16^{3}$) to evaluate the scalability, energy efficiency, and area efficiency of the EN-T architecture (Fig.3(c)). The benchmark object for testing is the PE composed of Synopsys DesignWare standard process library multiplier. The EN-T architecture utilizes two encoders (MBE and Our Encoder) with register outputs, and all test on 500MHz. The accuracy of multipliers in the PE is INT8 (Due to the prevalent use of INT8 in existing AI accelerators and edge devices, where the mantissa part of floating-point numbers often involves low-bit-width unsigned multiplication (BF16(UINT7), TF32(UINT10)), we adopt INT8 as the multiplication bit width for testing TCUs.), and the accumulator width is $16+\log_{2}S$. $S$ represents the array size.
Figure 9. SoC Normalized Energy Fraction under Baseline-based TCU
Figure 10. Single Frame Energy Consumption Comparison for SoC Inference Based on EN-T
@@ 4.3. Performance of EN-T architecture
In the tests of energy efficiency and area efficiency (Fig.7(a)(b)), we mainly compare the scalability of individual TCUs with different computational scales under EN-T architecture. Due to the square relationship between the number of removed encoders and the array size, increasing the computational scale to a certain extent will result in higher performance. As shown in Fig.7(a)(b), when the computational scale of our encoder-based EN-T architecture expands from 256GOPS to 1TOPS and 4TOPS, the average area efficiency improves from 8.7% to 12.2% and 11.0%, and the energy efficiency improves from 13.0% to 17.5% and 15.5%, respectively. Among them, the 1D/2D Array achieves a 20.2% increase in area efficiency and a 20.5% increase in energy efficiency compared to the baseline at 1TOPS. This is due to the specific characteristics of the multiplier-adder architecture itself (with no PEs, multipliers and multiplicands are not pipelined to the adder tree). In this case, the performance improvement of the EN-T architecture is the highest.
4.4. Benchmark on SoC
In this section, we will evaluate the performance of the EN-T architecture on System-on-Chip (SoC) level for neural network inference (ResNet34, ResNet50, ResNet101, Inception_V3, DenseNet121, DenseNet161, Vgg13, Vgg19). We use a basic NPU architecture and run on 500MHz, as shown in the Fig.8, which includes three levels of storage. The SoC contains two levels of on-chip storage: a 256KB Global Buffer and 64KB Activation and Weight Buffers. In the readout of the Weight Buffer, we have added 32 Encoders to convert weights into encoded numbers for computation in the TCUs, with the Encoder module using register output. The Controller module includes control for reading and writing to SRAM and contains an img2col module for preprocessing the convolution operations. We use five types of TCU architectures: 2D Matrix, 1D/2D Array, two types of Systolic Array (WS and OS), and 3D Cube, with an array size of 32$\times$32. The 3D Cube configuration consists of two $8^{3}$ arrays, with a unified computational scale of 1024GOPS. Externally to the TCUs, there is a SIMD Vector Processing Engine, internally using 32 ALUs with TF32 precision for quantization, pooling, scalar addition, and activation functions. The detailed hardware parameters of each module are shown in Table.2.
Figure 11. Energy Consumption Reduction Ratio of SoC
We conducted an energy consumption analysis for single-frame image (1,3,224,224) inference on the entire SoC, starting with an analysis of the energy proportion normalized for the SoC based on the Baseline TCU, as shown in Fig.9. We decomposed the on-chip energy consumption into the read and write energy of SRAM, as well as the energy generated by the computing engines (TCU and SIMD Vector Processing Engine). For neural network computations, the energy consumed by the computing engines accounts for a high proportion of the on-chip energy, ranging from 80-94%. This is due to the high data reuse rate inherent in convolution and matrix multiplication. For instance, in a Systolic Array, each read and write operation to SRAM allows data to be transmitted and undergo multiply-accumulate operations across all rows and columns of the array, making the computing array, rather than on-chip memory access, the primary source of power consumption. However, for some lightweight networks like DenseNet or Mobilenet that use a higher proportion of memory-intensive depthwise separable convolutions, the proportion of power consumption from memory access increases, but still does not exceed 25% (as shown in Fig.9(c)). Therefore, reducing the power consumption of the TCU is beneficial for lowering the system.
We replaced the Baseline TCU with the EN-T architecture and extracted the encoder tests for all the adders to analyze the energy consumption before and after the replacement. The energy consumption of the SoC is shown in Fig.10, and the energy reduction ratio is shown in Fig.11. For the 2D Matrix architecture, the energy consumption can be reduced by 15.1%-15.9% across different networks; for the Systolic Array (OS) architecture, can be reduced by 11.3%-12.8%; for the Systolic Array (WS) architecture, can be reduced by 10.2%-11.7%; for the 1D/2D Array architecture, can be reduced by 14.0%-16.0%; for the 3D Cube architecture, can be reduced by 5.0%-6.0%. Among these, applying the EN-T architecture under the 3D Cube scenario yields lower benefits compared to other architectures. This is due to the 3D structure requiring more encoders for the same computational power. For example, a 32$\times$32 two-dimensional array requires 32 encoders, saving 992 encoders, while to achieve 1024 GOPS of computational power with a 3D Cube, two $8^{3}$ arrays are needed, requiring 128 encoders and saving 896 encoders. Therefore, the energy saving benefits are not as significant as those of other 2D architectures.
Figure 12. Area Efficiency of the SoC
In terms of area efficiency, we compared the area efficiency improvement ratios of the Baseline as show in Fig.12, as well as the individual TCU and SoC. Due to the addition of on-chip SRAM, Controller, and SIMD Vector Processing Engine, as shown in Table.2, the area proportion of the on-chip SRAM is basically equivalent to that of the computing modules. Therefore, from the perspective of SoC, the area benefits brought by applying the EN-T architecture are relatively low. The main advantage is that it can reduce the inference power consumption by 10%-16%.
5. Conclusion
This paper explores new avenues in addressing the computational reuse issues in traditional TCUs by designing a matrix multiplication unit with an EN-T architecture. This method boasts high versatility and can be directly applied to existing tensor computing engines. From the perspective of computational encoding, we propose a method for encoding data with lower bit widths within the multiplier and uncover new opportunities for reducing energy consumption from the perspective of the multiplier’s internals. Our method combines the multiplier and adder calculation with matrix multiplication from a more fine-grained perspective and designs a low-bit width data coding method based on the carry chain. Overall, our scheme can inspire other researchers to reconsider the design of tensor computing units.